OpenStudy (anonymous):

this years, john played in 10 baseball games. In these games, he had hit the ball 2,3,0,1,3,2,4,0,2 and 3 times. In the first 10 games he plays next year, john wants to increase his average (mean) hits per game by 0.5. What is the total number of hits john needs over the first 10 games next year to achieve his goal?

OpenStudy (anonymous):

Total number of hits in first 10 games= 2+3+0+1+3+2+4+0+2+3 = 20 Current mean hits per game = total number of hits/number of games = 20/10 = 2 He wants to increase his mean hits up to 2.5 per game so, \[(20+x)/10 =2.5\] \[20+x=10\times2.5\] \[20+x=25\] \[x=5\] He needs to increase his total hits per 10 games by 5 So the total hits needed per game for next year = 20+5 =25 Hope that helps